OpenStudy (anonymous):

Find the limit of (lnx)^(1/x) as x approaches to infinity.

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(\ln x\right)^{1/x}\] \[\exp\left(\ln\lim_{x\to\infty}\left(\ln x\right)^{1/x}\right)\] \[\exp\left(\lim_{x\to\infty}\ln \left(\ln x\right)^{1/x}\right)\] \[\exp\left(\lim_{x\to\infty}\frac{\ln \left(\ln x\right)}{x}\right)\] A good candidate for L'hopital's rule